# Learning Category-level Last-meter Navigation from RGB Demonstrations of a Single-instance

### v1

@@ I INTRODUCTION
For robots to integrate effectively into human environments and perform useful tasks, they must be capable of both moving through the space and interacting with surrounding objects. This integration of navigation and manipulation, commonly referred to as mobile manipulation, is a fundamental capability for autonomous assistive robots. Crucially, manipulation policies operate reliably only when the robot’s position and orientation fall within the distribution covered by their demonstration data [1]. Existing navigation approaches, however, are designed around much coarser requirements: most global navigation benchmarks define success as stopping within roughly 1 meter of the target location [2, 3, 4, 5, 6, 7, 8]. Even the widely adopted navigation evaluation protocol from Anderson et al. [9] similarly specifies success as reaching within twice the agent’s body width, typically about 1 meter from the target location. As a result, navigation systems often fail to position the robot precisely enough for the manipulation system to follow and succeed, leading to frequent failures during task execution [10]. This mismatch reveals a critical gap between these phases of mobile manipulation.
@@
To achieve centimeter-level positioning precision, navigation systems often rely on strong prior knowledge of the environment, such as high-resolution maps or neural radiance fields (NeRFs) [11, 12, 5], or require additional sensing modalities such as depth cameras, LiDAR, or odometry [3, 4, 6, 7]. These requirements for prior knowledge and additional sensing modalities limit scalability in real-world deployment: environmental priors are costly to obtain and often brittle in dynamic or movable-target settings, while extra sensors add hardware cost and system complexity. In contrast, RGB cameras are inexpensive, already standard on most commercial and research robots, and often sufficient for many perception tasks, making RGB-only policies an appealing direction.
Recent navigation research therefore emphasizes RGB-based, end-to-end learning methods, either via imitation learning or reinforcement learning [10, 2], and vision-language foundation models that leverage large-scale pretraining for semantic navigation [4]. However, these methods still adopt coarse success metrics (e.g., one-meter thresholds), leaving their precision inadequate for downstream manipulation tasks that demand centimeter-level accuracy.
@@ II-A RGB-only Learning-based Global Navigation
Recent end-to-end learning approaches have enabled robots to navigate complex and visually diverse environments using only RGB observations. PoliFormer [2] presents a transformer-based on-policy reinforcement learning framework that scales efficiently to long-horizon navigation and achieves state-of-the-art results in challenging visual scenes. In a related approach, NoMaD [13] employs goal-conditioned diffusion policies that jointly model exploration and navigation, producing robust trajectories without explicit maps or depth input. While many other navigation systems leverage additional sensing modalities such as depth, LiDAR, or pre-built maps to simplify localization and planning [3, 14, 11, 5, 12, 6, 4, 7], PoliFormer and NoMaD demonstrate that RGB-only visual inputs combined with high-capacity learning architectures can still produce reliable global navigation policies. However, they typically focus on reaching the general vicinity of the goal, rather than achieving the fine-grained alignment necessary for manipulation. Our work addresses this limitation by explicitly focusing on the last-meter phase, where centimeter-level precision is essential for interaction with target objects.
@@ II-B Positioning Robot Base for Mobile Manipulation
Accurate robot base positioning is essential for reliable manipulation performance. Mobi-$\pi$ [1] clearly demonstrates that manipulation policies fail when the robot’s base pose falls outside the distribution of the training data, underscoring the importance of precise positioning. Similarly, optimization-based studies such as MoMa-Pos [15] and  [16] consistently emphasize that the choice of base location directly affects reachability, manipulability, and overall task success. Together, these findings confirm that base placement is a key determinant of manipulation success in mobile manipulation systems. Integrated frameworks such as SPOC [10] further highlight this dependency: despite combining navigation and manipulation within a unified policy, their reported failures often stem from slight misalignments between the gripper and the target object, reinforcing the critical need for accurate last-meter navigation.
@@ II-C Last-meter Navigation
Several recent works have begun addressing the last-meter navigation stage that directly precedes manipulation. Aim My Robot [17] introduces a precision local navigation policy that guides the robot toward arbitrary target objects with centimeter-level accuracy, using depth and LiDAR sensing for geometric reasoning. MoTo [18] extends this direction with an interaction-aware navigation policy that generates feasible docking points via object-centric reasoning. However, MoTo relies on depth sensing and an offline pre-scanning phase to construct a global scene graph, which prevents online adaptation in dynamic environments. While these methods successfully target the last-meter stage, their reliance on heavy sensing modalities and static global maps restricts general deployment. In contrast, our approach attains comparable precision using only RGB input, enabling broader applicability across diverse robots and settings.
@@ III-C Architecture
Recent advances in vision foundation models provide powerful pretrained encoders that map images into semantically rich embeddings [19, 20, 21, 22]. Our system leverages these pretrained representations to reduce the amount of task-specific training data, especially when there is no simulation environment available. The proposed architecture consists of three modules: (1) a vision encoder, (2) a segmentation module, and (3) an action decoder (Fig. 2).
@@ III-C1 Vision Encoder
We adopt DINOv2 [19, 20] as the vision encoder to extract robust, semantically aware image features from both the current and goal observations.
@@ III-C2 Segmentation Module
To highlight the target object, we employ a two-stage text-driven segmentation process. OwlV2 [23] detects the object specified by the language prompt and produces a bounding box, which is then refined by SAM2 [24] to generate a segmentation mask. The resulting bounding box and mask are passed downstream to crop and filter the visual embeddings for action prediction.
@@ V-A Baseline
The first baseline, DinoTxtAttention, follows a canonical multimodal attention pipeline. In this base line, DINOv2 is used as the vision encoder to extract embeddings from both the current and goal observations. A pretrained text encoder [25] maps the language prompt to a text embedding that matches the dimensionality of the DINOv2 embeddings. The text embedding then attends to both the goal and current visual embeddings. The resulting attended goal and current embeddings are passed through a cross-attention layer and a feedforward action head to predict the discrete actions.
@@ V-D How does the best policy perform on unseen instances in their own environment?
Figure 9: Ten unseen scenarios used for evaluating generalization.

### v2

@@ I INTRODUCTION
For robots to integrate effectively into human environments and perform useful tasks, they must be capable of both moving through the space and interacting with surrounding objects. This integration of navigation and manipulation, commonly referred to as mobile manipulation, is a fundamental capability for autonomous assistive robots. Crucially, manipulation policies operate reliably only when the robot’s position and orientation fall within the distribution covered by their demonstration data [undef]. Existing navigation approaches, however, are designed around much coarser requirements: most global navigation benchmarks define success as stopping within roughly 1 meter of the target location [undefa, undefb, undefc, undefd, undefe, undeff, undefg]. Even the widely adopted navigation evaluation protocol from Anderson et al. [undefh] similarly specifies success as reaching within twice the agent’s body width, typically about 1 meter from the target location. As a result, navigation systems often fail to position the robot precisely enough for the manipulation system to follow and succeed, leading to frequent failures during task execution [undefi]. This mismatch reveals a critical gap between these phases of mobile manipulation.
@@
To achieve centimeter-level positioning precision, navigation systems often rely on strong prior knowledge of the environment, such as high-resolution maps or neural radiance fields (NeRFs) [undefj, undefk, undefd], or require additional sensing modalities such as depth cameras, LiDAR, or odometry [undefb, undefc, undefe, undeff]. These requirements for prior knowledge and additional sensing modalities limit scalability in real-world deployment: environmental priors are costly to obtain and often brittle in dynamic or movable-target settings, while extra sensors add hardware cost and system complexity. In contrast, RGB cameras are inexpensive, already standard on most commercial and research robots, and often sufficient for many perception tasks, making RGB-only policies an appealing direction.
Recent navigation research therefore emphasizes RGB-based, end-to-end learning methods, either via imitation learning or reinforcement learning [undefi, undefa], and vision-language foundation models that leverage large-scale pretraining for semantic navigation [undefc]. However, these methods still adopt coarse success metrics (e.g., one-meter thresholds), leaving their precision inadequate for downstream manipulation tasks that demand centimeter-level accuracy.
@@ II-A RGB-only Learning-based Global Navigation
Recent end-to-end learning approaches have enabled robots to navigate complex and visually diverse environments using only RGB observations. PoliFormer [undefa] presents a transformer-based on-policy reinforcement learning framework that scales efficiently to long-horizon navigation and achieves state-of-the-art results in challenging visual scenes. In a related approach, NoMaD [undefl] employs goal-conditioned diffusion policies that jointly model exploration and navigation, producing robust trajectories without explicit maps or depth input. While many other navigation systems leverage additional sensing modalities such as depth, LiDAR, or pre-built maps to simplify localization and planning [undefb, undefm, undefj, undefd, undefk, undefe, undefc, undeff], PoliFormer and NoMaD demonstrate that RGB-only visual inputs combined with high-capacity learning architectures can still produce reliable global navigation policies. However, they typically focus on reaching the general vicinity of the goal, rather than achieving the fine-grained alignment necessary for manipulation. Our work addresses this limitation by explicitly focusing on the last-meter phase, where centimeter-level precision is essential for interaction with target objects.
@@ II-B Positioning Robot Base for Mobile Manipulation
Accurate robot base positioning is essential for reliable manipulation performance. Mobi-$\pi$ [undef] clearly demonstrates that manipulation policies fail when the robot’s base pose falls outside the distribution of the training data, underscoring the importance of precise positioning. Similarly, optimization-based studies such as MoMa-Pos [undefn] and  [undefo] consistently emphasize that the choice of base location directly affects reachability, manipulability, and overall task success. Together, these findings confirm that base placement is a key determinant of manipulation success in mobile manipulation systems. Integrated frameworks such as SPOC [undefi] further highlight this dependency: despite combining navigation and manipulation within a unified policy, their reported failures often stem from slight misalignments between the gripper and the target object, reinforcing the critical need for accurate last-meter navigation.
@@ II-C Last-meter Navigation
Several recent works have begun addressing the last-meter navigation stage that directly precedes manipulation. Aim My Robot [undefp] introduces a precision local navigation policy that guides the robot toward arbitrary target objects with centimeter-level accuracy, using depth and LiDAR sensing for geometric reasoning. MoTo [undefq] extends this direction with an interaction-aware navigation policy that generates feasible docking points via object-centric reasoning. However, MoTo relies on depth sensing and an offline pre-scanning phase to construct a global scene graph, which prevents online adaptation in dynamic environments. While these methods successfully target the last-meter stage, their reliance on heavy sensing modalities and static global maps restricts general deployment. In contrast, our approach attains comparable precision using only RGB input, enabling broader applicability across diverse robots and settings.
@@ III-C Architecture
Recent advances in vision foundation models provide powerful pretrained encoders that map images into semantically rich embeddings [undefr, undefs, undeft, undefu]. Our system leverages these pretrained representations to reduce the amount of task-specific training data, especially when there is no simulation environment available. The proposed architecture consists of three modules: (1) a vision encoder, (2) a segmentation module, and (3) an action decoder (Fig. 2).
@@ III-C1 Vision Encoder
We adopt DINOv2 [undefr, undefs] as the vision encoder to extract robust, semantically aware image features from both the current and goal observations.
@@ III-C2 Segmentation Module
To highlight the target object, we employ a two-stage text-driven segmentation process. OwlV2 [undefv] detects the object specified by the language prompt and produces a bounding box, which is then refined by SAM2 [undefw] to generate a segmentation mask. The resulting bounding box and mask are passed downstream to crop and filter the visual embeddings for action prediction.
@@ V-A Baseline
The first baseline, DinoTxtAttention, follows a canonical multimodal attention pipeline. In this base line, DINOv2 is used as the vision encoder to extract embeddings from both the current and goal observations. A pretrained text encoder [undefx] maps the language prompt to a text embedding that matches the dimensionality of the DINOv2 embeddings. The text embedding then attends to both the goal and current visual embeddings. The resulting attended goal and current embeddings are passed through a cross-attention layer and a feedforward action head to predict the discrete actions.
@@ V-D How does the best policy perform on unseen instances in their own environment?
Figure 9: Ten unseen scenarios used for evaluating generalization.